You decide to visit Santa Claus at the north pole to put in a good word about your splendid behavior throughout the year. While there, you notice that the elf Sneezy, when hanging from a rope, produces a tension of 425 N in the rope.

If Sneezy hangs from a similar rope while delivering presents at the earth's equator, what will the tension in it be? (Recall that the earth is rotating about an axis through its north and south poles.) Consult Appendix F in the textbook.
Answer #1

here,

mass of earth , M = 5.97 * 10^24 kg

the gravity at equator , g1 = 9.78 m/s^2

the gravity at poles , g2 = 9.832 m/s^2

the tension in the string at the equator , T1 = 425 N

the tension in the rope at North Pole , T2 = T1 * g2 /g1

T2 = 425 * 9.832 /9.78 N

T2 = 427.3 N